Kiara baked 30 oatmeal cookies and 48 chocolate chip cookies to package in plastic containers for her teacher friends at school. She wants to divide the cookies into identical containers so that each container has the same number of each kind of cookie. If she wants each container to have the greatest number of cookies possible, how many plastic containers does she need? *
1
Expert's answer
2019-12-17T10:31:58-0500

Firstly find the greatest common divisor of 30 and 48: 30 / 6 = 5; 48 / 6 = 8, so the greatest common divisor is 6, it means that Kiara needs 6 plastic containers which include 30 / 6 = 5 oatmeal cookies and 48 / 6 = 8 chip cookies.

Answer: 6 containers.